Use iteration, either forward or backward substitution, to solve the recurrence relation an=an−1−1 for any positive integer n, with initial condition, a0=1. Use mathematical induction to prove the solution you find is correct.


1
Expert's answer
2021-11-08T19:47:23-0500
"d=a_n-a_{n-1}=-1"

Arithmetic progression


"a, a+d, a+2d,..."

"a_n=a+d(n-1)"

"a_n=a-n+1"

Let "P(n)" be the proposition that "a_n=a-n+1" for any positive integer "n."

Basis Step:

"P(1)" is true, because

"a_1=a-1+1=a"

Inductive Step:

Assume that "P(k)" holds for an arbitrary positive integer "k." That is, we assume that


"a_k=a-k+1"

Under this assumption, it must be shown that "P(k + 1)" is true, namely, that


"a_{k+1}=a-(k+1)+1"

is also true.


"a_{k+1}=a_k\u22121"

Substitute


"a_{k+1}=a-k+1\u22121=a-(k+1)+1"

"a_{k+1}=a-(k+1)+1"

This last equation shows that "P(k + 1)" is true under the assumption that "P(k)" is true. This completes the inductive step

We have completed the basis step and the inductive step, so by mathematical induction we know that "P(n)" is true for all positive integers "n." That is, we have proven that


"a_n=a-n+1"

for all positive integers "n."